OpenStudy (anonymous):

Eight Question Of The Day (8. Q.O.T.D.)

OpenStudy (anonymous):

Are you happy now, @Hero?

OpenStudy (anonymous):

@anteater

OpenStudy (anteater):

Hello

OpenStudy (anteater):

Too many windows open!

OpenStudy (anteater):

Oh, ok ... it says to look for a Pythagorean triple. And they show a base triangle of 5-12-13, which must be a right triangle, since 5^2 + 12^2 = 13^2 (Pythagorean Theorem)

OpenStudy (anteater):

So you know you have that right triangle as your base. So 1/2 base times height for that thing would be (1/2)(5)(12) or 30

OpenStudy (anteater):

So the base of your pyramid is 30

OpenStudy (anteater):

And so 1/3 times base times height 1/3 times 30 times 9 = 90 m^3

OpenStudy (anteater):

And ... I think maybe you already got it answered, so I just showed up to verify it!

OpenStudy (anteater):

So, in case I missed you, that should be the answer! Off to do errands now. Hope you have a good afternoon!!!